a sled is pulled by a rope 25 degrees above horizontal. the sled has a mass of ? and the coefficent of friction is ? . the force applied to the rope is? n calculate the sleds acceleration....

enter in your own numbers into the questions marks but please tellme how i would go about solving this...i have a test tomorrow
1
Expert's answer
2012-01-20T07:22:37-0500
A sled is pulled by a rope α=25° above horizontal. the sled has a mass of M and the coefficent of friction is K. The force applied to the rope is F. Calculate the sleds acceleration.

The projection of the pulling forse at horizontal direction is

Fh = F*cos(α).

Now let's use the second Newton's law:

M*a = (1-K)*Fh ==> a = (1-K)*Fh/M = (1-K)*F*cos(α)/M.